40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Tue Jan 22 20:26:18 2019, awade, DailyProgress, TempCtrl, Running relay controller overnight on refcavs 
    Reply  Wed Jan 23 11:28:53 2019, awade, DailyProgress, TempCtrl, Correcting BN after running relay controller overnight on refcavs 
       Reply  Thu Jan 24 14:24:10 2019, awade, DailyProgress, TempCtrl, Custom implementation of PIDLocker script for finer control of intergration on cav heaters PIDLocker_cavcontrol.py.zip
          Reply  Fri Jan 25 10:46:24 2019, awade, DailyProgress, TempCtrl, PID locking of cavity heaters to BN PID_locking_GlitchAndRunAway_2019-01-25_at_10.50.19_AM.png
             Reply  Fri Jan 25 15:45:48 2019, anchal, DailyProgress, TempCtrl, PID locking of cavity heaters to BN 
Message ID: 2289     Entry time: Tue Jan 22 20:26:18 2019     Reply to this: 2290
Author: awade 
Type: DailyProgress 
Category: TempCtrl 
Subject: Running relay controller overnight on refcavs 

Beat note is still not under control. Its drifting at 10 kHz/min, too much for a lower range marconi measurement.

We would like to bring the frequency down from ~140 MHz  to something closer to 27.34 MHz  of Koji's resonant BN detector.  I'm going to try to use the relay auto locker feature of the PIDLocker_beta.py script again to get the BN to converge on some steady value (see PSL:2142). In the past the relay method was able to actually get a convergence to a BN value but failed to give good values as there was a thing with the cycle limit converging to zero in the case of these really slow BN thermal controls.  I suspect this was due to a relay step that was too small, the asymmetry of the heater actuation and/or large time lag caused issues.

For now I'll leave the relay on overnight to see if we are able to get a good convergence with a relay step of 0.1 W about mean heater power of 0.77320 W and common power offset of 0.48710 W.  The setpoint was left at 100 MHz, we will step it down to the 27.43 MHz gradually so as not to overshoot the BN past 0 Hz and flip the direction of the frequency.

The service running the PIDLocker_beta.py script for the cavity slow heater controls was stopped and was then manually restarted in a tmux session on ws1 using the following command:

> python PIDLocker_beta.py PIDConfig_CAVHeater.ini --autotune -d 0.1 -o 0.7732 -t 86400 -e

This should run the relay mode of the auto locker for 24 hours (86400 s) about a mean of 0.7732 W and a relay step of 0.1 W (bigger than in previous tests).  We'll see how it goes and reassess tomorrow morning.

ELOG V3.1.3-